# ScarFinder: a detector of optimal scar trajectories in quantum many-body dynamics

### v1

@@ III Spin-1 XY Model
One of the simplest and most studied models that contain analytically exact QMBSs is the spin-1 XY chain [17, 80]:
@@
where $\hat{S}_{j}^{\alpha}$ are the standard spin-1 operators with $\alpha=x,y,z$ on site $j$, $h$ is the external magnetic field in the $z$-direction, and $\hat{V}$ is a perturbation term that breaks integrability and any symmetries that are not essential for QMBS states. In previous works [17, 80], it was shown that the model (7) hosts two families of QMBS eigenstates, known as the Type-1 and Type-2 scar towers. The two types are distinguished by the initial state they overlap with; as we detail below, the Type-I tower overlaps with a product state, while the other, Type-II, tower has overlap with an MPS state of bond dimension $\chi=2$. From the perspective of ScarFinder, the algorithm performs similarly in the two cases, hence in the remainder of this section we focus on Type-1 scar tower, while the results for Type-2 are relegated to Appendix C.
@@
Here, |+ket|$\rangle$| + ⟩ and |ket|$\rangle$| - ⟩ denote the highest and lowest weight spin-1 basis states, respectively.
The scar eigenstates $|\psi_{n}\rangle$ in Eq. (9) can be coherently combined to form special initial states that exhibit persistent, nonthermal dynamics under unitary evolution [17]:
@@
We refer to this quantity as the logarithmic fidelity. Since the scar states form a continuous submanifold $\mathcal{S}$, we introduce a fidelity measure – the “scar fidelity” – to quantify how closely a given state $\psi$ aligns with $\mathcal{S}$:
@@
We remark that during iterative applications of ScarFinder, the state’s energy may drift due to bond truncation. Although the energy density of the scar state in Eq. (11) covers the interval $[-h,+h]$, uncontrolled energy drift could complicate precise trajectory selection. This is why it is essential to enforce energy targeting, as previously explained in Sec. II.2.
Figure 3: Convergence behavior of ScarFinder initialized from random states at target energy $E_{\text{target}}=0$, with $\Delta t=0.2$. Energy conservation is explicitly enforced during the algorithm’s iterations. (a) For the $\hat{V}=\hat{V}_{1}$ perturbation in Eq. (8), among 100 independent trials, only 5 successfully converge. (b) For the $\hat{V}=\hat{V}_{1}^{\prime}$ perturbation in Eq. (15), 9 out of 10 trials converge successfully. (c)-(d): Entanglement entropy of eigenstates for (c) $\hat{V}=\hat{V}_{1}$ and (d) $\hat{V}=\hat{V}_{1}^{\prime}$. Improved convergence in (b) correlates with a clearer separation between scar and thermal states in the entanglement entropy spectrum. Data in panels (a)-(b) is obtained using iTEBD with a maximal bond dimension $\chi=16$ and evolution step $dt=0.01$. Panels (c)-(d) are exact diagonalization results for a finite system $L=10$, combining momentum sectors $k=0$ and $k=\pi$.
To illustrate sensitivity to initial conditions, we run the energy-corrected ScarFinder algorithm with randomly initialized states in Fig. 3(a). For the perturbation $\hat{V}_{1}$, the success rate is low, highlighting significant initial-state dependence. The reason for this is the simplicity of $\hat{V}_{1}$ perturbation, which leaves some residual structure in the non-scar eigenstates. To achieve a fully generic scenario and to rigorously test the theoretical assumptions underlying the ScarFinder algorithm, we next consider a more general perturbation
@@
where $\hat{\lambda}^{(i)}$ are Gell-Mann matrices. This perturbation can be derived from the quantum inverse method, as detailed in Appendix A. As shown in Fig. 3(b), the generic $\hat{V}_{1}^{\prime}$ perturbation substantially improves convergence rates.
@@
Finally, for the spin-1 XY model, in addition to the Type-1 QMBS states discussed above, there exists a distinct Type-2 scar tower [17, 80]. This Type-2 scar tower features a weakly-entangled scar trajectory described by MPS. In Appendix C we demonstrate that the ScarFinder successfully obtains the Type-2 trajectory, with similar convergence criteria as in the Type-1 case illustrated above.
@@ IV PXP model
In this section we turn our attention to the experimentally relevant example of QMBSs in the PXP model, which is natively realized in Rydberg atom arrays [14, 61] and has also been engineered using tilt potential in a Bose-Hubbard optical lattice [23]. The PXP model [51, 52, 15] represents a one-dimensional spin-1/2 chain with the Hamiltonian:
@@
[…]
The PXP model is known to be chaotic [15], while at the same time it also hosts a small number of non-thermal QMBS eigenstates which are evenly distributed in energy and possess anomalously low entanglement entropy [15, 70, 81, 57]. These QMBS eigenstates have high overlap with the $|\mathbb{Z}_{2}\rangle\equiv|{\downarrow}{\uparrow}{\downarrow}{\uparrow}{\cdots}\rangle$ state, leading to a suppressed growth of entanglement entropy and periodic revivals in local observables when the system is quenched from that state. The resulting scar dynamics were shown to have an elegant semiclassical explanation in terms of a variational manifold $\mathcal{M}$ spanned by $\chi=2$ MPS states [53]. Equivalently, the scar dynamics can be interpreted as precession of a collective spin in an approximate su(2) RSGA picture [70, 54, 50, 59]. The latter approach has also allowed to construct deformations of the PXP model which lead to nearly perfect revivals from the $|\mathbb{Z}_{2}\rangle$ initial state [54, 56, 50, 59].
The PXP revival dynamics is understood to be a many-body analog of scar phenomena in stadium billiards, which are quantum remnants of classical, unstable periodic orbits [82, 83, 84, 85]. Nevertheless, some important questions remain open. For example, the QMBS subspaces beyond those associated with $|\mathbb{Z}_{2}\rangle$ are poorly understood, a simple example being the reviving $|\mathbb{Z}_{3}\rangle\equiv|{\downarrow}{\downarrow}{\uparrow}{\downarrow}{\downarrow}{\uparrow}{\cdots}\rangle$ state with a repeated 3-site unit cell [70]. This leads to the question: can one do better than $|\mathbb{Z}_{2}\rangle$ and $|\mathbb{Z}_{3}\rangle$ states, i.e., can one construct low-entangled initial states that have more pronounced revivals than these product states? Specifically, for the undeformed PXP model, is it possible to find initial states that exhibit nearly perfect revival dynamics and almost no entanglement growth in the thermodynamic limit? In the remainder of this section, we demonstrate that ScarFinder algorithm can provide useful insights into these questions.
@@ IV.1 Two-site unit cell
Figure 4: (a)-(b) Dynamics of the optimal scar initial state for the PXP Hamiltonian with a two-site unit cell and various bond dimensions $\chi$. (a) Entanglement entropy growth. Compared to the $|\mathbb{Z}_{2}\rangle$ state (black line), $\chi=2$ already shows significantly reduced entanglement growth, while $\chi=8,12$ dynamics are nearly flat, indicating an extremely stable scar. (b) Logarithmic fidelity dynamics for the optimal state at $\chi=12$, shows nearly perfect revivals with no visible decay. (c)-(f) Overlap between the optimized scar state $|\psi(\chi)\rangle$ and the eigenstates of the PXP Hamiltonian for different bond dimensions: (c) $\chi=2$, (d) $\chi=4$, (e) $\chi=8$, and (f) $\chi=12$. As $\chi$ increases, a well-defined band of scarred eigenstates becomes more distinct from the thermal bulk, with increasing spectral isolation. Data in panels (a)-(b) are obtained by iTEBD, while (c)-(f) are exact diagonalization results for system size $L=24$ in $k=0$ momentum sector. Since iTEBD simulations in (a)-(b) use a 4-site unit cell, the momentum $k=0$ in exact diagonalization data corresponds to translation by 4 sites.
In order to understand the mechanism for the dynamics observed in Fig. 4(a)-(b), it is useful to inspect the eigenstate composition of the optimal scar state. As shown in Fig. 4(c)-(f), the states belonging to the scar band generally split into two subsets as $\chi$ is increased. The subset closer to the middle of spectrum develops an enhanced overlap with the optimal scar state, while the overlaps of the other subset diminish and approach those of typical eigenstates in the bulk of the spectrum. Thus, an increase in $\chi$ results in a clearer separation between some of the scarred eigenstates and the other eigenstates. Importantly, we find that these high-overlap eigenstates coincide with those observed in the $|\mathbb{Z}_{2}\rangle$ case. Thus, an increase in $\chi$ effectively tunes the coefficients of these scar eigenstates and picks out a superposition that approximates better and better the optimal scar state.
IV.2 Three-site unit cell and energy conservation
Beyond the well-known two-site unit cell dynamics, the PXP model also supports approximate periodic revivals on a three-site unit cell, most notably when the system is prepared in the $|\mathbb{Z}_{3}\rangle\equiv\left|\uparrow\downarrow\downarrow\uparrow\downarrow\downarrow\cdots\right\rangle$ state [15, 70]. These revivals, however, are much weaker than those associated with the $|\mathbb{Z}_{2}\rangle$ initial state. Ref. [71] constructed an MPS state with $\chi=2$ that showed improved revival fidelity compared to the simple $|\mathbb{Z}_{3}\rangle$ product state. We now apply the ScarFinder algorithm to the three-site unit cell in order to systematically explore how well the $|\mathbb{Z}_{3}\rangle$ revivals can be improved by increasing the amount of entanglement in the initial state.
@@ IV.2 Three-site unit cell and energy conservation
The procedure is identical to the two-site case, except for adapting the iMPS ansatz to a three-site periodic structure. However, during the random sampling of initial states, we often find states whose dynamics are governed by the same set of eigenstates identified in the two-site unit cell setting. Thus, the $\mathbb{Z}_{2}$ orbit acts as an attractor for the three-site dynamics. The appearance of $\mathbb{Z}_{2}$-type oscillations in the three-site unit cell setting can be understood as follows: the $|\mathbb{Z}_{2}\rangle$ initial state has nonzero overlap with eigenstates in the momentum $k=0$ sector defined under one-site translation symmetry. This $k=0$ sector is fully contained within the $k=0$ sector defined under three-site translation symmetry. As a result, we naturally expect to observe remnants of the $\mathbb{Z}_{2}$-type scar dynamics even when working with a three-site unit cell. To isolate genuine three-site scar trajectories, we perform a post-selection based on the oscillation frequency, allowing us to filter out spurious results dominated by enhanced overlaps with these residual $\mathbb{Z}_{2}$ modes. These are easily distinguished by their revival period: the intrinsic three-site scar dynamics feature a shorter oscillation period compared to those from the two-site case [70].
@@
Figure 5: (a)-(b) Dynamics of the optimized scar initial states for the PXP Hamiltonian with a 3-site unit cell and different bond dimensions $\chi$. (a) Late-time entanglement entropy growth ($t\in[15,30]$). Increasing $\chi$ systematically suppresses the entanglement growth and stabilizes periodic dynamics. The “*-state” refers to the MPS-optimized initial state from Ref. [71], previously identified as having better revivals compared to the $|\mathbb{Z}_{3}\rangle$ product state. (b) Late-time ($t\in[15,30]$) logarithmic fidelity for the optimized initial state at $\chi=12$, computed from the dominant eigenvalue of the MPS transfer matrix. The fidelity displays persistent revivals, highlighting the stability of the identified scar trajectory. (c)-(f) Overlaps between energy eigenstates and the optimized initial scar state for bond dimensions $\chi=2$ (c), $\chi=4$ (d), $\chi=8$ (e), and $\chi=12$ (f). With increasing $\chi$, distinct eigenstate towers clearly emerge from the thermal continuum, signifying improved scar dynamics. Data in panels (c)-(f) is obtained by exact diagonalization for a system size $L=24$ in the $k=0$ momentum sector w.r.t. 3-site translations.
@@
The resulting optimal trajectories in a 3-site unit cell are shown in Fig. 5(a). For comparison, the so-called “*-state” introduced in Ref. [71] was constructed using a fixed MPS ansatz with bond dimension $\chi=2$. We observe that with increasing bond dimension, the entanglement growth of the optimized scar states is further reduced, and the fidelity revivals become more pronounced. For $\chi=12$ [Fig. 5(b)], the fidelity revival is nearly perfect, although not as high as in the two-site unit cell case.
@@
In Fig. 5(c)-(f), we also plot the overlaps of the initial scar states with the energy eigenstates. Interestingly, unlike the two-site unit cell case where the overlap distribution is more dispersed, the three-site unit cell dynamics produce discrete towers of states with approximately equal energy spacing, highlighting a distinct and more structured scar subspace.

### v3

@@ III Spin-1 XY Model
One of the simplest and most studied models that contain analytically exact QMBSs is the spin-1 XY chain [Schecter2019, Chattopadhyay2020]:
@@
where $\hat{S}_{j}^{\alpha}$ are the standard spin-1 operators with $\alpha=x,y,z$ on site $j$, $h$ is the external magnetic field in the $z$-direction, and $\hat{V}$ is a perturbation term that breaks integrability and any symmetries that are not essential for QMBS states. In previous works [Schecter2019, Chattopadhyay2020], it was shown that the model (7) hosts two families of QMBS eigenstates, known as the Type-1 and Type-2 scar towers. The two types are distinguished by the initial state they overlap with; as we detail below, the Type-I tower overlaps with a product state, while the other, Type-II, tower has overlap with an MPS state of bond dimension $\chi=2$. From the perspective of ScarFinder, the algorithm performs similarly in the two cases, hence in the remainder of this section we focus on Type-1 scar tower, while the results for Type-2 are relegated to Appendix C.
@@
Here, $|+\rangle$ and $|$-$\rangle$ denote the highest and lowest weight spin-1 basis states, respectively.
The scar eigenstates $|\psi_{n}\rangle$ in Eq. (9) can be coherently combined to form special initial states that exhibit persistent, nonthermal dynamics under unitary evolution [Schecter2019]:
@@
We refer to this quantity as the logarithmic fidelity. Note that the fidelity revival implies recurrence of all local observables. Since the scar states form a continuous submanifold $\mathcal{S}$, we introduce a fidelity measure – the “scar fidelity” – to quantify how closely a given state $\psi$ aligns with $\mathcal{S}$:
@@
We remark that during iterative applications of ScarFinder, the state’s energy may drift due to bond truncation. Although the energy density of the scar state in Eq. (11) covers the interval [h,+h][-h,+h], uncontrolled energy drift could complicate precise trajectory selection. This is why it is essential to enforce energy targeting, as previously explained in Sec. II.2.
Figure 3: Convergence behavior of ScarFinder initialized from random states at target energy $E_{\text{target}}=0$, with $\Delta t=0.2$. Energy conservation is explicitly enforced during the algorithm’s iterations. (a) For the $\hat{V}=\hat{V}_{1}$ perturbation in Eq. (8), among 100 independent trials, only 5 successfully converge. (b) For the $\hat{V}=\hat{V}_{1}^{\prime}$ perturbation in Eq. (15), 9 out of 10 trials converge successfully. (c),(d): Entanglement entropy of eigenstates for (c) $\hat{V}=\hat{V}_{1}$ and (d) $\hat{V}=\hat{V}_{1}^{\prime}$. Improved convergence in (b) correlates with a clearer separation between scar and thermal states in the entanglement entropy spectrum. Data in panels (a),(b) are obtained using iTEBD with a maximal bond dimension $\chi=16$ and evolution step $dt=0.01$. Panels (c)-(d) are exact diagonalization results for a finite system $L=10$, combining momentum sectors $k=0$ and $k=\pi$.
To illustrate sensitivity to initial conditions, we run the energy-corrected ScarFinder algorithm with randomly initialized states in Fig. 3(a). For the perturbation $\hat{V}_{1}$, the success rate is low, highlighting significant initial-state dependence. The reason for this is the simplicity of $\hat{V}_{1}$ perturbation, which leaves some residual structure in the non-scar eigenstates. To test the assumptions underlying the ScarFinder algorithm, we next consider a more general perturbation
@@
where $\hat{\lambda}^{(i)}$ are Gell-Mann matrices. This perturbation can be derived from the quantum inverse method, as detailed in Appendix A. While neither perturbation is fully generic due to this constraint, the second one includes a broader range of local terms and thus breaks more of the model’s residual structure. As shown in Fig. 3(b), the $\hat{V}_{1}^{\prime}$ perturbation indeed substantially improves the convergence rate.
@@
Finally, for the spin-1 XY model, in addition to the Type-1 QMBS states discussed above, there exists a distinct Type-2 scar tower [Schecter2019, Chattopadhyay2020]. This Type-2 scar tower features a weakly-entangled scar trajectory described by MPS. In Appendix C we demonstrate that the ScarFinder successfully obtains the Type-2 trajectory, with similar convergence criteria as in the Type-1 case illustrated above.
@@ IV PXP model
In this section we turn our attention to the experimentally relevant example of QMBSs in the PXP model, which is natively realized in Rydberg atom arrays [bernien2017probing, bluvstein2021controlling] and has also been engineered using tilt potential in a Bose-Hubbard optical lattice [Su2023]. The PXP model [FendleySachdev, Lesanovsky2012, turner2018weak] represents a one-dimensional spin-1/2 chain with the Hamiltonian:
@@
where $\Omega=1$ is the Rabi frequency, $\hat{\sigma}_{j}^{x}$ is the standard Pauli-$x$ operator on site $j$, and $\hat{P}_{j}=(1-\hat{\sigma}_{j}^{z})/2$ is the projector on the local $|{\downarrow}\rangle$ state. The PXP model describes constrained flipping of atoms in the Rydberg blockade regime [Browaeys2020]: each atom can flip only if both of its neighbors are in the $|\downarrow\rangle$ state. Hence, neighboring excitations, such as $|{\cdots}{\uparrow}{\uparrow}{\cdots}\rangle$, are energetically forbidden, which imposes a constraint on the dynamics.
The PXP model is known to be chaotic [turner2018weak], while at the same time it also hosts a small number of non-thermal QMBS eigenstates which are evenly distributed in energy and possess anomalously low entanglement entropy [turner2018weak, Turner2018PRB, IadecolaSchecterXu2019, lin2019exact]. These QMBS eigenstates have high overlap with the $|\mathbb{Z}_{2}\rangle\equiv|{\downarrow}{\uparrow}{\downarrow}{\uparrow}{\cdots}\rangle$ state, leading to a suppressed growth of entanglement entropy and periodic revivals in local observables when the system is quenched from that state. The resulting scar dynamics were shown to have an elegant semiclassical explanation in terms of a variational manifold $\mathcal{M}$ spanned by $\chi=2$ MPS states [Ho2019]. Equivalently, the scar dynamics can be interpreted as precession of a collective spin in an approximate su(2) RSGA picture [Turner2018PRB, Choi2019, Bull2020, Omiya2022]. The latter approach has also allowed to construct deformations of the PXP model which lead to nearly perfect revivals from the $|\mathbb{Z}_{2}\rangle$ initial state [Choi2019, Khemani2019, Bull2020, Omiya2022].
The PXP revival dynamics is understood to be a many-body analog of scar phenomena in stadium billiards, which are quantum remnants of classical, unstable periodic orbits [Heller84, HellerLesHouches, Bogomolny1988, Berry1989Scars]. Nevertheless, some important questions remain open. For example, the QMBS subspaces beyond those associated with $|\mathbb{Z}_{2}\rangle$ are poorly understood, a simple example being the reviving $|\mathbb{Z}_{3}\rangle\equiv|{\downarrow}{\downarrow}{\uparrow}{\downarrow}{\downarrow}{\uparrow}$$\cdots$$\rangle$ state with a repeated 3-site unit cell [Turner2018PRB]. This leads to the question: can one do better than $|\mathbb{Z}_{2}\rangle$ and $|\mathbb{Z}_{3}\rangle$ states, i.e., can one construct low-entangled initial states that have more pronounced revivals than these product states? Specifically, for the undeformed PXP model, is it possible to find initial states that exhibit nearly perfect revival dynamics and almost no entanglement growth in the thermodynamic limit? In the remainder of this section, we demonstrate that ScarFinder algorithm can provide useful insights into these questions.
@@ IV.1 Two-site unit cell
Figure 4: (a)-(c) Dynamics of the optimal scar initial state for the PXP Hamiltonian with a two-site unit cell and various bond dimensions $\chi$. (a) Entanglement entropy growth. Compared to the $|\mathbb{Z}_{2}\rangle$ state (black line), $\chi=2$ already shows significantly reduced entanglement growth, while $\chi=8,12$ dynamics are nearly flat, indicating an extremely stable scar. (b) Logarithmic fidelity dynamics for the optimal state at $\chi=12$, shows nearly perfect revivals with no visible decay. (c) Expectation value of the nearest-neighbor observable $\langle Z_{i}Z_{i+1}\rangle$ for $\chi=12$, showing persistent oscillations throughout the evolution. (d)-(g) Overlap between the optimized scar state $|\psi(\chi)\rangle$ and the eigenstates of the PXP Hamiltonian for different bond dimensions: (d) $\chi=2$, (e) $\chi=4$, (f) $\chi=8$, and (g) $\chi=12$. As $\chi$ increases, a well-defined band of scarred eigenstates becomes more distinct from the thermal bulk, with increasing spectral isolation. Data in panels (a)-(c) are obtained by iTEBD, while (d)-(g) are exact diagonalization results for system size $L=24$ in $k=0$ momentum sector. Since iTEBD simulations in (a)-(c) use a 4-site unit cell, the momentum $k=0$ in exact diagonalization data corresponds to translation by 4 sites.
We run ScarFinder on $\hat{H}_{\mathrm{PXP}}^{\prime}$ by sampling 100 random initial states in $\mathcal{M}$ and selecting one that exhibits the lowest entanglement entropy at time t=30t=30.111We use the iTEBD algorithm to simulate the time evolution. Since local terms in the PXP Hamiltonian act on three neighboring sites, a four-site unit cell is required to implement the Trotter decomposition. Nevertheless, the resulting dynamics largely preserve the underlying two-site translational symmetry. The resulting optimal dynamics are shown in Fig. 4(a)-(c). Notably, we find that even with a relatively small bond dimension $\chi=8$, the algorithm produces an initial state that shows minimal entanglement growth and remarkably robust fidelity revivals—signatures of scar dynamics that persist over long times ($t\lesssim 30$) in the thermodynamic limit. We remark that the results presented here are based on minimal entanglement entropy criterion; however, they are not the only scar solutions identified by ScarFinder. Specifically, for $\chi=3$ (and also $\chi=4$), we also find the exact $E=0$ MPS eigenstate of the PXP Hamiltonian reported in Ref. [lin2019exact].
In order to understand the mechanism for the dynamics observed in Fig. 4(a)-(c), it is useful to inspect the eigenstate composition of the optimal scar state. As shown in Fig. 4(d)-(g), the states belonging to the scar band generally split into two subsets as $\chi$ is increased. The subset closer to the middle of spectrum develops an enhanced overlap with the optimal scar state, while the overlaps of the other subset diminish and approach those of typical eigenstates in the bulk of the spectrum. Thus, an increase in $\chi$ results in a clearer separation between some of the scarred eigenstates and the other eigenstates. Importantly, we find that these high-overlap eigenstates coincide with those observed in the $|\mathbb{Z}_{2}\rangle$ case. Thus, an increase in $\chi$ effectively tunes the coefficients of these scar eigenstates and picks out a superposition that approximates better and better the optimal scar state. A qualitatively similar enhancement of overlaps near the middle of the spectrum was observed in Ref. [PhysRevB.102.195114], albeit in a different setting where the PXP Hamiltonian was perturbed to enhance the scar dynamics.
IV.2 Three-site unit cell and energy conservation
@@
The procedure is identical to the two-site case, except for adapting the iMPS ansatz to a three-site periodic structure. However, during the random sampling of initial states, we often find states whose dynamics are governed by the same set of eigenstates identified in the two-site unit cell setting. Thus, the $\mathbb{Z}_{2}$ orbit acts as an attractor for the three-site dynamics. The appearance of $\mathbb{Z}_{2}$-type oscillations in the three-site unit cell setting can be understood as follows: the $|\mathbb{Z}_{2}\rangle$ initial state has nonzero overlap with eigenstates in the momentum $k=0$ sector defined under one-site translation symmetry. This $k=0$ sector is fully contained within the $k=0$ sector defined under three-site translation symmetry. As a result, we naturally expect to observe remnants of the $\mathbb{Z}_{2}$-type scar dynamics even when working with a three-site unit cell. To isolate genuine three-site scar trajectories, we perform a post-selection based on the oscillation frequency, allowing us to filter out spurious results dominated by enhanced overlaps with these residual $\mathbb{Z}_{2}$ modes. These are easily distinguished by their revival period: the intrinsic three-site scar dynamics feature a shorter oscillation period compared to those from the two-site case [Turner2018PRB].
@@
Figure 5: (a)-(c) Dynamics of the optimized scar initial states for the PXP Hamiltonian with a 3-site unit cell and different bond dimensions $\chi$. (a) Late-time entanglement entropy growth ($t\in[15,30]$). Increasing $\chi$ systematically suppresses the entanglement growth and stabilizes periodic dynamics. The “*-state” refers to the MPS-optimized initial state from Ref. [Michailidis2020], previously shown to have better revivals compared to the $|\mathbb{Z}_{3}\rangle$ product state. (b) Late-time ($t\in[15,30]$) logarithmic fidelity for the optimized initial state at $\chi=12$, computed from the dominant eigenvalue of the MPS transfer matrix. The fidelity displays persistent revivals, highlighting the stability of the identified scar trajectory. (c) Time evolution of the nearest-neighbor observable $\langle Z_{i}Z_{i+1}\rangle$ for $\chi=12$. (d)-(g) Overlaps between energy eigenstates and the optimized initial scar state for bond dimensions $\chi=2$ (d), $\chi=4$ (e), $\chi=8$ (f), and $\chi=12$ (g). With increasing $\chi$, distinct eigenstate towers clearly emerge from the thermal continuum, signifying improved scar dynamics. Data in panels (d)-(g) is obtained by exact diagonalization for a system size $L=24$ in the $k=0$ momentum sector w.r.t. 3-site translations.
@@
The resulting optimal trajectories in a 3-site unit cell are shown in Fig. 5(a). For comparison, the so-called “*-state” introduced in Ref. [Michailidis2020] was constructed using a fixed MPS ansatz with bond dimension $\chi=2$. We observe that with increasing bond dimension, the entanglement growth of the optimized scar states is further reduced, and the fidelity revivals become more pronounced. For $\chi=12$ [Fig. 5(b)], the fidelity revival is nearly perfect, although not as high as in the two-site unit cell case. In addition, the improved scar dynamics is further supported by the oscillatory behavior of the nearest-neighbor observable $\langle Z_{i}Z_{i+1}\rangle$, shown in Fig. 5(c).
@@
In Fig. 5(d)-(g), we also plot the overlaps of the initial scar states with the energy eigenstates. Interestingly, unlike the two-site unit cell case where the overlap distribution is more dispersed, the three-site unit cell dynamics produce discrete towers of states with approximately equal energy spacing, highlighting a distinct and more structured scar subspace.